OpenStudy (anonymous):

what is the distance in centimeters from goal line to goal line on a standard football field?

OpenStudy (wolf1728):

That is 100 yards which is 300 feet or 3,600 inches (300*12). There are 2.54 centimeters per inch. Can you make the final calculation?

OpenStudy (anonymous):

no i'm confused

OpenStudy (theeric):

Are you working on learning how to convert between different distances?

OpenStudy (anonymous):

yess I really don't know how to convert though

OpenStudy (anonymous):

Multiply 3,600 inches by 2.54. Then you'll have your answer.

OpenStudy (anonymous):

9,144 centimeters. That's your answer. You could also just Google it. LOL

OpenStudy (anonymous):

Thankkk u

OpenStudy (anonymous):

No problem. :)

OpenStudy (anonymous):

Please mark my answer as best and give me mY MEDAAAAAALLLLL!!!!!!

OpenStudy (anonymous):

lol jk

OpenStudy (anonymous):

did u get the medal?

OpenStudy (theeric):

I see you have your answer, but I'll tell you about how to convert most things... You use what's called a "conversion factor." It's equal to \(1\), so you can multiply it by any amount without changing the amount. How do you get it so it works nicely? Well, \(1\ yard=3\ feet\), so you divide both sides by one of the sides. I'll divide by \(1\ yard\). Then \(1= \dfrac{3\ feet}{1\ yard}\) So, you have \(100\ yards\). \(100\ yards \times 1\\=100\ yards\times\dfrac{3\ feet}{1\ yard}\\=100\ \cancel{yards}\times\dfrac{3\ feet}{1\ \cancel{yard}}\\=300\ feet\)

OpenStudy (anonymous):

ahhh thanks that helped. I wish I could medal this

OpenStudy (theeric):

Haha, no problem! Glad I could help! :)

OpenStudy (theeric):

If you wanted to continue the problem in this way, you'd use \(1\ inch=2.54\ cm\) to get to your new conversion factor. You can make \(\dfrac{cm}{inch}\) or \(\dfrac{inch}{cm}\) depending on how you do it! You want inches to cancel out, so you'll have inches on the bottom.

OpenStudy (anonymous):

Nice job, Eric!

OpenStudy (theeric):

Thank you, @micahpkay

OpenStudy (theeric):

!

OpenStudy (anonymous):

Your welcome. :)

OpenStudy (anonymous):

And yes, I did get my medal, @ebailey258. Thanks so much!